OpenStudy (anonymous):

Find all polar coordinates of point P where P= (6,- pi/ 5).

OpenStudy (anonymous):

ok....is there a picture you can post to help us with this question

OpenStudy (anonymous):

nope

OpenStudy (mathmale):

The coordinates of a point can be expressed in either cartesian or polar coordinates. In this case there's no need to worry about cartesian coordinates. r = the radius = the distance of the given point from the origin = 6 theta = the angle measured from the positive x-axis = - pi/5 You could convert -pi/5 to degrees if you wish. Graph the point (6, -pi/5) From your graph we can begin to answer the question you've posted.

OpenStudy (mathmale):

\[Hint:-\frac{ \pi }{ 5 }\frac{ 180~ \deg }{ \pi }=?~\deg\]

OpenStudy (anonymous):

What do I do afterwards?

OpenStudy (mathmale):

Have you graphed (6,-pi/5)? Please share your work ... use the Draw facility (below).

OpenStudy (anonymous):

 i think that is really wrong